OpenStudy (anonymous):

Does anybody know how to do this? If the population of India was 650 million a decade ago and is now 790 million people and continues to grow exponentially at the same rate,what will the population be in 5 years?

OpenStudy (amistre64):

790 = 650(1+r)^10

OpenStudy (anonymous):

can you please help me step by step.please because im having a hard time with this problem

OpenStudy (amistre64):

A=P(1+r)^t A/P = (1+r)^t (A/P)^(1/t) = 1+r (A/P)^(1/t) - 1 = r Ao = 650(1+((790/650)^(1/10)) - 1)^15 maybe

OpenStudy (anonymous):

thankyou for helping me ;)

OpenStudy (amistre64):

do you have options to check against?

OpenStudy (anonymous):

no i dont

OpenStudy (amistre64):

hmm, then im not sure what setup to use; i spose we could try a continual setup and see if its close to what this one is

OpenStudy (amistre64):

A=Pe^(rt) A/P=e^(rt) ln(A/P) = rt ln(A/P)/t = r A = 650e^(15*ln(790/650)/10) hmm, that works out remarkably well too

OpenStudy (amistre64):

both give me 870.93; so id say 871 just so we dont have any .93 people walking about

OpenStudy (amistre64):

or tp be more exacter; 870,930,000 might suffice

OpenStudy (anonymous):

thank you for your help.

OpenStudy (amistre64):

yep, hope it was readable